OpenStudy (ltrout):

[Medal] The half-life of a material is 42 days. An initial amount of the material has a mass of 49 kg. Write an exponential function that models the decay of this material.

OpenStudy (ltrout):

I'm thinking it's going to be y = 49(1/2)^_x Idk what to put for the _ before the x.

OpenStudy (ltrout):

@Qwertty123 Quick help?

OpenStudy (qwertty123):

Okay... The half-life of a certain radioactive material is 42 days. An initial amount of the material has a mass of 49 kg. Write an exponential function that models the decay of this material. : A = Ao*2^(-t/h), where: A = amt after t time Ao = initial amt (t=0) t = time of decay

OpenStudy (qwertty123):

h = half life of substance : A = 49*2^(-t/42) : : Find how much radioactive material remains after 8 days. A = 49*2^(-8/42) A = ?

OpenStudy (ltrout):

It would be about 42.940 Kg. So the equation would be \[y = 49\frac{ 1 }{ 2 }^{\frac{ 1 } { 42 } x}\]?

OpenStudy (qwertty123):

yes it would be..

OpenStudy (ltrout):

Alright. Thank you :)

OpenStudy (qwertty123):

A = 42.9395 ~ 42.940, to the nearest thousandth.

OpenStudy (qwertty123):

You welcome..